OpenStudy (anonymous):

Intrinsic value according to the DDM = D1/(k-g), where the P/E is1/(k-g). In english, what is the p/e formula telling us? Thanks!

OpenStudy (anonymous):

The P/E formula is telling us that the firm is expected to pay out all of its earnings as dividends. P=D1/(k-g) ->P/E=(D1/E)*(1/(k-g)) = 1/(k-g) if D1=E

OpenStudy (anonymous):

P/E=(1-b)/(k-g), IT SAYS THAT THE FIRM WILL KEEP PAYING THIS CASH TO INVESTORS TILL PERPETUITY. b=RETENTION RATIO 1-b IS DIVIDEND PAYOUT.